OpenStudy (anonymous):

Essay: Show all work. Find the following quotient: 25x^2-100y^2/5x+10y

OpenStudy (anonymous):

Difference of 2 squares in the numerator, can you factor those?

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

look familiar?

OpenStudy (anonymous):

yes but i never know how to plug the numbers in

OpenStudy (anonymous):

what is\[\sqrt{25x^2}\]

OpenStudy (anonymous):

5x^2

OpenStudy (anonymous):

close

OpenStudy (anonymous):

what is\[\sqrt{x^2}\]

OpenStudy (anonymous):

x^1.41

OpenStudy (anonymous):

\[\sqrt{x^2}=x\]

OpenStudy (anonymous):

can you explain that rick? and then retrice natz the original question?

OpenStudy (anonymous):

weird bug, ask natz*

OpenStudy (anonymous):

huh dgo

OpenStudy (anonymous):

i typed a s s k i think

OpenStudy (anonymous):

anyhow I am wondering what you are taking away from all the answers youve been getting. are you learning how to do this stuff?

OpenStudy (anonymous):

yes i have learned a lot from this website i just have a hard time grasping it .It takes me to learn it mre than once before i got it all the way

OpenStudy (anonymous):

your answer is 5x-10y use difference of squares to simplify into: \[(5x+10y)(5x-10y)\div(5x+10y)\] (5x+10y) crosses out from the equation because \[(5x+10y)\div(5x+10y)=1\] so you are left with 5x-10y

OpenStudy (anonymous):

same for me natz haha ive learned a lot from this website, just takes me so long

OpenStudy (anonymous):

it is just math is my worst subject

OpenStudy (anonymous):

addrigonzalez is that the whole problem

OpenStudy (anonymous):

do you understand adri's answer?

OpenStudy (anonymous):

it should be, you are looking for a quotient, right? so if 6 is divided by 3, the quotient is 2

OpenStudy (anonymous):

yes they cross out eachother

OpenStudy (anonymous):

i love math, we are on derivatives in college and i just dont understand my foreign teacher or whay we even need derivatives in the first place

OpenStudy (anonymous):

derivatives are unnessecary